OpenStudy (anonymous):

For f(x) = x2 − 5x + 3, find and simplify the following. 4f(x)

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 5x + 3 4*f(x) = 4*(x^2 - 5x + 3) ... Multiply both sides by 4. 4*f(x) = ???

OpenStudy (anonymous):

4x^2-20x+12

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

thanks:)

jimthompson5910 (jim_thompson5910):

you're welcome